OpenStudy (anonymous):

write a function rule relating x and y x=1,2,3 y=1,8,27 i cant see the pattern these are my weak point ehh

OpenStudy (anonymous):

\[y=x^3\] 1^3=1 2^3=8 3^3=27

OpenStudy (anonymous):

sorry im a tab bit confused the letters with what y = is kinda hard to see is it x cubed =y?

OpenStudy (amistre64):

yes

OpenStudy (amistre64):

the ascii text beloow it kinda gives it away :)

OpenStudy (anonymous):

ohh okay